OpenStudy (anonymous):

studying for my final, please check my work! find the quadratic function that satisfies f(1)=0, f(-1)=14 and f(2)=2 i got 3a-7b+4c

OpenStudy (asnaseer):

you seem to have calculated the right values for a, b, and c for a quadratic equation of the form: \(ax^2+bx+c=0\) but have written the answer down incorrectly.

OpenStudy (asnaseer):

sorry I meant: \(ax^2+bx+c=f(x)\)

OpenStudy (anonymous):

how would i right it?

OpenStudy (asnaseer):

so what you need to do is substitute for a, b and c into this equation.

OpenStudy (asnaseer):

you found: a=3 b=-7 c=4

OpenStudy (anonymous):

oh so would it be 9a-7b+4c

OpenStudy (asnaseer):

no - you need to substitute the values for a, b and into this equation:\[f(x)=ax^2+bx+c\]so you should end up with an equation that just has numbers and the variable 'x' within it (\(x^2\) and \(x\) that it).

OpenStudy (asnaseer):

e.g. you know a=3, so: \(f(x)=3x^2+bx+c\) now substitute in the values for b and c as well.

OpenStudy (anonymous):

oh so just \[3x ^{2}-7x+4\]

OpenStudy (asnaseer):

thats right - well done! :)

OpenStudy (anonymous):

thank you!

OpenStudy (asnaseer):

yw :)